An ambulance driving 20.0 m/s emits a sound wave with a frequency of 500 Hz. As it drives away from a hospital, which of the following is the only possible frequency heard by the medical staff who are standing at the entrance? 


1
Expert's answer
2022-03-06T18:06:14-0500

The Doppler's effect equation says

"f=f_0\frac{c}{c+v_s}"

"f=500*\frac{340}{340+20}=472\:\rm Hz"